OpenStudy (anonymous):

2/3x^2 - x = 1/3 help

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

substract 1/3 from both sides. Then solve as a quadratic equation.

OpenStudy (saifoo.khan):

x = 1.78 y = -0.28

OpenStudy (anonymous):

I need in soulution set

OpenStudy (saifoo.khan):

{1.78, -0.28}

jhonyy9 (jhonyy9):

sorry but if you have thought this like : (2/3x^2) -x =1/3 ---- so if yes than you will get 2-3x^3 =x^2 --- x^2 =y --- than we get 2-3y^2=y --- 3y^2 +y-2=0 -1+/- sqrt(1+24) -1+/- 5 y_1= 4/6 = 2/3 y_1,_2= ---------------- = -------- = y_2 = -6/6 = -1 6 6 y_1 =2/3 ---- x^2 = 2/3 --- x_1,_2 =+/- sqrt(2/3) y_2 = -1 ---- x^2 = -1 --- x_3 = i --- because i^2 = -1